OpenStudy (anonymous):

Is This TRUE or FALSE??The slope-intercept form of the equation of the line with slope 0 passing through the point (-2,7) is y = 7.

OpenStudy (anonymous):

slope of 0 is a horizontal line, so true because the y-coordinate is 7

OpenStudy (anonymous):

Give the equation of the line that contains the points (-8, -4) and (-7,-4)

OpenStudy (anonymous):

is this the correct answery= 3/16x -72

OpenStudy (anonymous):

no, because you should have a horizontal line here with a 0 slope because -4-(-4)=0

OpenStudy (anonymous):

y=-4 So this can be the answer?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Find the equation of a line with slope -4 and y-intercept (0,1)

OpenStudy (anonymous):

y=-4x+1 this is the answer i got is it correct?

OpenStudy (anonymous):

use point-slope form of line y-y1=m(x-x1) y-1=-4(x-0) y-1=-4x y=-4x+1

OpenStudy (anonymous):

Sarah makes decorative wreaths and sells them on eBay. She recently discovered that if she spends no money on advertising her wreaths, she can sell 100 wreaths in a year. However, if she spends $200 on advertising, her sales go up to 150 wreaths in a year. There is a linear relationship between the money spent on advertising and the number of wreaths sold. If x represents the amount spent on advertising and y represents the number of wreaths that Sarah sells, write an equation for y in terms of x

OpenStudy (anonymous):

y = 1/4x + 150 this is the answer i got